OpenStudy (anonymous):

Suppose all the dimensions of a right rectangular prism are multiplied by 5. a. what happens to the Surface Area? b. what happens to its volume?

OpenStudy (anonymous):

I just need my answers checked.. a. 150 times bigger b.125 times bigger. Thanks!

OpenStudy (experimentx):

surface area get multiplied by 5x5 times while volume gets multiplied by 5x5x5 times

OpenStudy (anonymous):

A. surface area of rectangular prism is 6xL^2, if it's 5L then 6x(5L)^2= 150 times bigger B. volume= L^3, if 5L then (5L)^3= 125 times bigger. you're correct